Please show steps to solve this problem x-2/3=5/6
Answer by kritikapd(4)

x-2/3=5/6
LCM of X-2/3 is 3
3X-2/3 = 5/6
cross mutiplying both sides
6(3x-2)= 3(5)
18x-12 = 15
18x = 15+12 = 27
x = 27/18 = 3/2
Ans: x=3/2
Verify
3/2-2/3= 5/6
5/6 = 5/6 Equation satisfied